# Understanding Biases in ChatGPT-based Recommender Systems: Provider Fairness, Temporal Stability, and Recency

### v1

@@ 5.3. Diversity and Provider Fairness
Gini Coefficient HHI Entropy Gini Coefficient HHI Entropy
@@
5.4. Temporal Freshness Bias in Movie Recommendations
In this section, we study the recency and freshness of the movie recommendations provided by various GPT-based and CF models. The investigation focuses on whether the models demonstrate a bias toward recommending newer or older films. By examining the mean and median release years of the movies suggested by each model, we can gauge their tendency toward recency.
@@ 5.4. Temporal Freshness Bias in Movie Recommendations
Table 7 presents the statistics that inform our understanding of each model’s tendency towards recommending newer or older films. Notably, the GPT-based models ‘Surprise’ and ‘Diversify’ push the boundaries towards recommending more recent items, with median release years of 2006 and 2007. This indicates a clear orientation towards newer movies, possibly interpreting ‘surprise’ and ‘diversification’ in their temporal sense. Since GPT models do not possess inherent information about the novelty of content in terms of user interactions, they likely interpret these directives based on the freshness of the content itself. In contrast, the CF models tend to recommend older films, median release years consistently at 1995, reflecting a stronger inclination towards well-established, historically popular films.
This distinction suggests that GPT-based models might be leveraging their extensive training datasets to prioritize more recent films, which can be beneficial in scenarios where up-to-date recommendations are desired. The broader temporal range of recommendations from GPT-based models, as evidenced by the standard deviation in the year of movies recommended (Surprise: Std Year 10.62, Diversify: Std Year 11.76), further underscores their flexibility in catering to diverse temporal tastes.
@@
Answer to H1: Contrary to the hypothesis, the results reveals that GPT-based models, especially ‘Diversify’ (Median Year: 2007) and ‘Surprise’ (Median Year: 2006), tend to recommend significantly newer movies than CF models. CF models such as BPR-MF (Median Year: 1995) and Pop (Median Year: 1995) generally suggest older films, as reflected in the lower mean and median years. The larger standard deviation in years for GPT-based models like ‘Surprise’ (Std Year: 10.62) also indicates a broader temporal range in their recommendations. This suggests that GPT-based models, perhaps due to their ability to capture more recent trends and discussions from diverse datasets, are capable of recommending more recent content than CF models, thus refuting the initial hypothesis.
5.5. Analysis of Genre Bias/Affinity in Recommender Systems
The goal of this section is to examine the genre biases present in movie recommendations provided by Collaborative Filtering (CF) models and GPT-based models. By comparing the prevalence of genre terms in the recommendations, we aim to understand if a particular model, or class of models, has a specific genre affinity (or bias) which could illuminate the performance differences between these models.
H1: We hypothesize that the differences in recommendation performance between CF and GPT-based models may be attributed to their respective genre preferences. CF models are expected to closely adhere to genre affinity in users’ profiles, while GPT models may exhibit a predominant recommendation of certain genres over others.
Upon analyzing the word clouds, it is evident that CF models such as BPR-MF, LightGCN, and VAE predominantly recommend genres such as “Action,” “Adventure,” “Thriller,” and “Sci-Fi.” These genres appear frequently and in larger font sizes, indicating their dominance in the recommendations made by these models.
In contrast, GPT-based models, represented in the lower word clouds, show a more varied genre distribution with “Comedy,” “Drama,” and “Romance” and even “Thriller” and “Adventure” being more prominent. These results are interesting and suggest that GPT-based RecLLMs may be less constrained by the users’ historical genre affinities and more exploratory in their recommendations, possibly leading to a broader but potentially less precise genre match to individual user profiles.
@@ 5.5. Analysis of Genre Bias/Affinity in Recommender Systems
Answer to H1: The analysis supports the hypothesis to a certain extent. While CF models tend to reinforce existing user preferences by frequently recommending popular genres such as “Action” and “Sci-Fi,” GPT-based models demonstrate a broader genre coverage in their recommendations. The presence of a more diverse set of genres like “Comedy,” “Drama,” and “Romance” in GPT-based recommendations suggests these models may be utilizing their extensive training on diverse data sources to introduce a wider variety of content. This could potentially lead to a discovery of new interests for users, though it may sometimes diverge from their established preferences. Such a tendency for broader genre representation by GPT-based models could be a contributing factor to their distinct performance characteristics compared to CF models.
5.6. Implicit vs. explicit scenario
Analyzing the impact of including ratings in the prompt shows noteworthy results. In our study, we initially avoided using ratings for the GPT models, aligning with the implicit setting of the Collaborative Filtering (CF) baseline. However, for completeness, we explored the effect of revealing user ratings for selected items.
@@
These results suggest that including ratings significantly enhances the recommendation system’s performance. It provides a more personalized and accurate reflection of user preferences, which is a key aspect in recommendation systems. While other GPT-based models showed varied performance, the inclusion of ratings seems to offer a more consistent and improved outcome. This insight could be valuable in refining other scenarios and models, potentially leading to better overall system performance.
@@ 6. Conclusion and Future Directions
This research presents a comprehensive analysis of Recommender Systems using Large Language Models (RecLLMs), specifically focusing on ChatGPT-based systems. The study emphasizes the nuanced capabilities and inherent biases of these systems, highlighting their contrasting behaviors compared to traditional CF in the domain of movie recommendations. Our experimental findings underscore the significant impact of ‘prompt design’ strategies on recommendation quality, touching upon aspects such as accuracy, provider fairness, diversity, temporal stability, genre dominance, and freshness (recency).
We discovered that role-based prompts (those using clear system roles such as you a Recommender, or you are a Fair recommender) enhance fairness and diversity, effectively mitigating popularity bias. Interestingly, GPT-based models, while not always paralleling the performance of CF baselines, exhibit a preference for recommending newer and more diverse movie genres.
Moreover, our experiments reveal that including ratings in prompts appears to improve the personalization aspect of recommendations. GPT-based models also have a distinct inclination towards recommending newer, particularly those released post-2000. We hypothesize that a key reason for certain discrepancies in recommendations quality is that the GPT model might suggest movies not included in the dataset. By limiting the GPT model to choose only from a predefined set of test items, we anticipate a notable improvement in both the accuracy and relevance of its recommendations. Additionally, transitioning to a zero-shot scenario could significantly enhance personalization results in this context. In genral, we think that it important to balance personalization power of CF models with the broader exploratory nature of GPT-based RecLLMs, to ensure a diverse and engaging user experience.
Key areas for future research include:
  1. (1)
    Refine Prompt Design Strategies and Dataset Constraints. This involves developing prompts that effectively balance user preferences with broader content exploration, while also constraining the system to select items only from within the dataset. This limitation is expected to improve the accuracy and relevance of recommendations.
  2. (2)
    Explore Few-shot Learning to Improve accuracy and Contextual Relevance. Few-shot learning involves exposing the model to a small number of examples from the target domain, which can enhance its ability to make more informed recommendations.
  3. (3)
    Integrate GPT with CF in a RAG framework. Aimed to create a hybrid approach that leverages the strengths of both methodologies. This combination could potentially lead to more robust and effective recommender systems.
  4. (4)
    Expand research to cross-domain applications and conduct long-term user engagement studies. This includes evaluating RecLLMs in various content domains (e.g., books, music, e-commerce) and studying user interaction over extended periods to gauge satisfaction, trust, and the evolving nature of user preferences.
By focusing on these key areas, future research can significantly advance the field of recommender systems, particularly in leveraging the capabilities of large language models like GPT in diverse and dynamic recommendation scenarios.

### v3

@@ 5.1.3. Item Fairness
[…]
@@
Cyan shows the best performing methods.
Green shows good performing methods (relative to others).
Yellow shows lower performance models.
Results and Discussion
We now address the above research hypotheses, using Table 8 to inform our discussion. We primarily use the Gini coefficient and entropy for this analysis, noting that HHI shows somewhat similar performance to the Gini coefficient. Overall, by examining the Gini indices and entropy values, we observe that two CF models, Item-KNN and NGCF, provide the best performances in terms of item fairness. Among the RecLLM models, the Genre-focused, Motivated Reasoning, and Chain-of-thought scenarios exhibit notable item fairness after the CF baselines.
Summarizing the best-performing models, we have: $ItemKNN/NGCF>Genre/Motivate/COT>rest$. One notable and interesting result from the table is that asking GPT to produce novel or diverse recommendations has little or no significant impact on improving its recommendability. The entropy generally increases for RecLLM models when switching to Fair recommenders, indicating higher novelty in recommendations. This pattern is not observed in the CF models, which show no change in their metrics. To address H4, we find that GPT-based models generally exhibit lower item fairness compared to classical CF models, as hypothesized. However, contrary to the expectation, asking the system to produce more diverse or additional recommendations did not significantly increase item fairness. For H5, integrating a “Fair Recommender” role in GPT-based systems did enhance recommendation fairness, as indicated by the increased entropy in the RecLLM models. This suggests a mitigation of item unfairness and an enhancement of recommendation diversity.
Answer to Hypotheses
We can answer to the above hypotheses as following:
  • H4: $\circ$ Partially Supported — While classical CF models such as Item-KNN and NGCF exhibit higher fairness compared to RecLLM models, the GPT-based models named ”Genre-focused” and ”Motivated Reasoning” showed improvements in fairness. This hypothesis is partially supported as diversification strategies and requests for more novel recommendations did not increase item fairness in RecLLM models.
  • H5: $\checkmark$ Supported — Implementing a ”Fair Recommender” role within GPT-based systems leads to noticeable improvements in fairness. This supports the hypothesis that transitioning from normal to fair recommender systems enhances fairness, demonstrating the controllability in RecLLM models, particularly in terms of diversity and novelty.
5.1.4. Analysis of Temporal Freshness and Genre Bias in Recommender Systems
In this unified analysis, we explore the tendencies of various GPT-based and CF models in terms of recency and genre preferences in movie recommendations. Our study aims to discern whether these models demonstrate a bias towards recommending newer or older films and if they exhibit specific genre preferences that could illuminate performance differences between model types.
Research Hypotheses
We provide the following statement of research hypotheses:
  • H6. We hypothesize that GPT-based models, given their extensive training on diverse and recent datasets, are more likely to recommend newer films compared to classical CF models, which may favor older, well-established films.
  • H7. We expect that GPT-based models will show a broader genre distribution in their recommendations due to their training on diverse data sources, compared to CF models, which may adhere more strictly to users’ historical genre preferences.
@@ 5.1.4. Analysis of Temporal Freshness and Genre Bias in Recommender Systems
Results and Discussions
Table 9 presents the statistics that inform our understanding of each model tendency towards recommending newer or older films. Notably, the GPT-based models ‘Surprise’ and ‘Diversify’ push the boundaries towards recommending more recent items, with median release years of 2006 and 2007. This indicates a clear orientation towards newer movies, possibly interpreting ‘surprise’ and ‘diversification’ in their temporal sense. Since GPT models do not possess inherent information about the novelty of content in terms of user interactions, they likely interpret these directives based on the freshness of the content itself. In contrast, the CF models tend to recommend older films, median release years consistently at 1995, reflecting a stronger inclination towards well-established, historically popular films. This distinction suggests that GPT-based models might be leveraging their extensive training datasets to prioritize more recent films, which can be beneficial in scenarios where up-to-date recommendations are desired. The broader temporal range of recommendations from GPT-based models, as evidenced by the standard deviation in the year of movies recommended (Surprise: Std Year 10.62, Diversify: Std Year 11.76), further underscores their flexibility in catering to diverse temporal tastes.
Upon analyzing the word clouds, it is evident that CF models such as BPR-MF, LightGCN, and VAE predominantly recommend genres such as “Action,” “Adventure,” “Thriller,” and “Sci-Fi.” These genres appear frequently and in larger font sizes, indicating their dominance in the recommendations made by these models. In contrast, GPT-based models, represented in the lower word clouds, show a more varied genre distribution with “Comedy,” “Drama,” and “Romance” and even “Thriller” and “Adventure” being more prominent. These results are interesting and suggest that GPT-based RecLLMs may be less constrained by the users’ historical genre affinities and more exploratory in their recommendations, possibly leading to a broader but potentially less precise genre match to individual user profiles.
@@
  • H6. $\checkmark$ Supported – GPT-based models, particularly Diversify and Surprise, tend to recommend significantly newer movies than CF models. This finding supports the hypothesis, confirming that GPT-based models are effective in capturing recent trends and offering more temporally relevant recommendations.
  • H7. $\circ$ Partially Supported – While CF models are found to predominantly recommend genres such as Action and Sci-Fi, GPT-based models demonstrate a broader genre distribution, including Comedy, Drama, and Romance. This diversity in GPT-based recommendations supports the hypothesis to a certain extent, indicating that these models are utilizing their extensive training on diverse datasets to introduce a wider variety of content.
Implicit vs. explicit scenario. Analyzing the impact of including ratings in the prompt shows noteworthy results. In our study, we initially avoided using ratings for the GPT models, aligning with the implicit setting of the Collaborative Filtering (CF) baseline. However, for completeness, we explored the effect of revealing user ratings for selected items.
@@
These results suggest that including ratings significantly enhances the recommendation performance of the system. It provides a more personalized and accurate reflection of user preferences, which is a key aspect in recommendation systems. While other GPT-based models showed varied performance, the inclusion of ratings seems to offer a more consistent and improved outcome. This insight could be valuable in refining other scenarios and models, potentially leading to better overall system performance.
5.2. Sequential Recommendations
This section focuses on the performance of RecLLMs in sequential recommendation task, measured by recommendation top-$K$ accuracy, item fairness, and coverage. Results are presented in Table 10, with each generative scenario (using ChatGPT) in the upper sections and CF baselines in the bottom section.
Hypotheses
To investigate the effectiveness of different recommendation strategies, we propose the following hypotheses:
  • H1. Incorporating sequential in-context learning (ICL) prompts, including ICL-1 and ICL-2, results in improved recommendation accuracy compared to zero-shot learning.
  • H2. Sequential in-context learning (ICL) prompts will result in better item fairness and coverage compared to zero-shot learning.
  • H3. A combination of improved accuracy, item fairness, and coverage will be observed in certain settings, highlighting the optimal recommendation strategy.
Discussion.
Our evaluation highlights several findings regarding the personalization of recommendations using RecLLM.
  • Regarding H1, the results indicate that recommendation accuracy, measured by Hit Rate (HR) and NDCG, varies with different learning strategies. For the MovieLens dataset, without demographic information, the HR for zero-shot learning is 0.179, which decreases slightly for ICL-1 (0.154) and ICL-2 (0.150). In the LastFM dataset, zero-shot learning HR is 0.204, which decreases to 0.154 for both ICL-1 and ICL-2.
    However, including demographic information, especially in the Age-Group category, improves accuracy in the movie domain. For instance, in the MovieLens dataset with age-group context, the HR for ICL-2 improves to 0.204, which is higher than zero-shot learning with such information (0.171) or zero-shot learning with no demographic information (0.179). Another clear trend is the impact of the Recency-Focused (Rec.-Freq) sampling in the context part on recommendation accuracy. When sensitive attributes are included, Rec-Freq and ICL-2 generally show better performance compared to the same age-category included in zero-shot learning. For example:
    • In the MovieLens dataset with age-group context, the HR for Rec-Freq ICL-2 is 0.204, compared to 0.171 for zero-shot learning.
    • In the MovieLens dataset with gender context, the HR for Rec-Freq ICL-2 is 0.175, compared to 0.158 for zero-shot learning.
    • In the LastFM dataset with age-group context, the HR for Rec-Freq ICL-2 is 0.204, compared to 0.263 for zero-shot learning.
    • In the LastFM dataset with gender context, the HR for Rec-Freq ICL-2 is 0.204, compared to 0.221 for zero-shot learning.
    In terms of NDCG, however, results generally favor zero-shot learning scenarios across different combinations of demographic information and sampling strategies in both datasets. While there is a slight improvement in NDCG for the Rec-Freq ICL-2 scenario in the LastFM dataset compared to other scenarios (zero-shot or ICL-1), the differences are marginal. For example:
    • In the LastFM dataset with age-group context, the NDCG for Rec-Freq ICL-2 is 0.60000, compared to 0.56250 for ICL-1, and 0.82500 for zero-shot learning.
    • In the LastFM dataset with gender context, the NDCG for Rec-Freq ICL-2 is 0.65625, compared to 0.59375 for ICL-1, and 0.64375 for zero-shot learning.
    These examples illustrate that incorporating sensitive attributes and employing the Rec-Freq sampling method in ICL-2 can enhance recommendation accuracy, although the improvements in NDCG are not as pronounced.
  • Regarding H2, item fairness and coverage show mixed results across different settings. For item fairness, as measured by Gini and Entropy, zero-shot learning and ICL strategies show slight differences. For instance, in the MovieLens dataset with no demographic info, Gini for zero-shot learning is 0.57739 (lower is better), which increases to 0.66303 for ICL-1 and 0.65835 for ICL-2, indicating worse fairness. Entropy, where higher values are better, decreases from 3.47952 (zero-shot) to 2.87493 (ICL-1) and 2.88530 (ICL-2), suggesting that ICL prompts might lead to slightly worse item fairness. Coverage remains consistently low across all scenarios in the movie domain, and zero-shot learning provides considerably better coverage performance. For example:
    • In the MovieLens dataset without demographic information, the coverage for zero-shot learning is 0.02621, compared to 0.01685 for ICL-1 and ICL-2.
    • In the MovieLens dataset with gender context, the coverage for zero-shot learning is 0.02845, compared to 0.01498 for ICL-2.
    In the music domain, results generally favor ICL-2. For example:
    • In the LastFM dataset with age-group context, the coverage for Rec-Freq ICL-2 is 0.00210, compared to 0.00198 for zero-shot learning.
    • In the LastFM dataset with gender context, the coverage for Rec-Freq ICL-2 is 0.00218, compared to 0.00211 for zero-shot learning.
    Notably, including demographic information generally enhances fairness metrics. For example, with Gender context in MovieLens, the Gini index improves, and Entropy increases, indicating fairer item distribution.
  • Regarding H3, a combination of improved accuracy, item fairness, and coverage is generally observed with the Recency-Focused approach in both zero-shot and ICL-2 scenarios. By visually examining the placements of highlighted values, we note that Recency-Focused zero-shot and ICL-2 tend to provide the best combined performance of accuracy, item fairness, and coverage. Additionally, revealing demographic information, particularly age-group (young/old) and gender, has a noticeable positive impact on the fairness and coverage of the methods, thereby enhancing the utility of the system from both consumer and producer perspective.
Table 10. Recommendation alignment between ($\mathcal{R}_{m}$,  $\mathcal{R}_{m}^{a}$) based on Item Similarity $\beta_{item}$
Attribute Type ICL MovieLens LastFM
Acc. Item Fairness Cov. Acc. Item Fairness Cov.
HR NDCG Gini Entropy - HR NDCG Gini Entropy -
No Info Freq. 0-shot 0.179 0.53750 0.57739 3.47952 0.02621 0.204 0.66875 0.24882 4.89089 0.00195
ICL-1 0.154 0.55000 0.66303 2.87493 0.01685 0.154 0.47500 0.24464 4.91536 0.00201
ICL-2 0.150 0.40000 0.65835 2.88530 0.01685 0.154 0.44375 0.25137 4.92419 0.00201
Rec.-Freq. 0-shot 0.208 0.65000 0.55075 3.66855 0.03033 0.225 0.65000 0.18229 5.07136 0.00220
ICL-1 0.162 0.54375 0.60812 3.08940 0.01760 0.183 0.53750 0.20217 4.99446 0.00206
ICL-2 0.175 0.56875 0.64473 2.92146 0.01610 0.208 0.58125 0.20439 5.00125 0.00205
Gender Freq. 0-shot 0.146 0.58750 0.56064 3.60813 0.02845 0.196 0.68750 0.21299 5.00305 0.00211
ICL-1 0.167 0.50625 0.67541 2.71790 0.01460 0.158 0.51875 0.26066 4.82855 0.00188
ICL-2 0.171 0.43125 0.66145 2.76958 0.01498 0.154 0.42500 0.22654 4.99862 0.00215
Rec.-Freq. 0-shot 0.158 0.55000 0.54779 3.68560 0.02958 0.221 0.64375 0.18254 5.08343 0.00223
ICL-1 0.154 0.46250 0.64715 2.99814 0.01909 0.208 0.59375 0.19480 5.01322 0.00209
ICL-2 0.175 0.50625 0.64839 2.73183 0.01310 0.204 0.65625 0.18622 5.06197 0.00218
Age-Group Freq. 0-shot 0.154 0.70000 0.54193 3.72305 0.03145 0.183 0.56875 0.24498 4.91475 0.00198
ICL-1 0.175 0.57500 0.67186 2.73694 0.01498 0.146 0.48750 0.24892 4.90102 0.00199
ICL-2 0.150 0.45625 0.64637 2.87851 0.01572 0.146 0.41250 0.23557 4.97022 0.00211
Rec.-Freq. 0-shot 0.171 0.66875 0.56433 3.63473 0.02995 0.263 0.82500 0.18816 5.07481 0.00223
ICL-1 0.158 0.46875 0.63580 2.92408 0.01610 0.188 0.56250 0.20575 4.96456 0.00200
ICL-2 0.204 0.60000 0.63038 2.93595 0.01498 0.204 0.60000 0.19985 5.01733 0.00210
Cyan shows the best performing methods.
Green shows good performing methods (relative to others).
Yellow shows lower performance models.
Answer to Hypotheses
To address the proposed hypotheses, we present the following findings:
  • H1. $\circ$ Partially Supported – The incorporation of sequential in-context learning prompts (ICL-1 and ICL-2) improves recommendation accuracy in certain contexts, particularly with demographic information (e.g., Age-Group in MovieLens), though not consistently across all settings.
  • H2. $\times$ Rejected – Sequential in-context learning prompts do not consistently lead to better item fairness and coverage compared to zero-shot learning. In some cases, they worsen item fairness, as indicated by higher Gini indices and lower Entropy values.
  • H3. $\checkmark$ Supported – Certain settings, particularly Recency-Focused ICL-2 with demographic information, show a balance between accuracy and item fairness, indicating potential optimal recommendation strategies.
5.3. Economic and Practical Implications
Integrating large language models (LLMs) like GPT-3.5-turbo into recommender systems involves several economic and practical considerations. Below, we discuss the inference costs, latency issues, and potential mitigation strategies.
Inference Costs
Using LLMs involves significant API costs. For example, the cost per API call can be calculated as:
Cost per call=(tokens per $\text{Cost per call}=\left(\frac{\text{tokens per call}}{1000}\right)\times\alpha$
where $\alpha$ is the cost per 1000 tokens. For a system with $n$ users and $m$ prompts, the total cost $C$ is:
$C=n\times m\times\text{Cost per call}$
For example, if an average recommendation call uses 1500 tokens, the cost per call is $0.03$. For 610 users with 7 prompts each, the total cost would be $128.10$. To mitigate these costs, we can explore strategies such as token optimization, batch processing, and selective use of LLMs. Our code implements these strategies by limiting the maximum tokens per response and grouping prompts to minimize the number of API calls.
Latency
Latency, the time taken to generate recommendations, affects user experience. Observed latencies ranged from a few to over ten seconds per call. Mitigation strategies include asynchronous processing to handle multiple requests concurrently, caching mechanisms to store common recommendations, and using faster, distilled versions of LLMs. Our code addresses latency by measuring and optimizing the time taken for each API call and incorporating mechanisms to batch and streamline requests. Additionally, practical considerations such as using cloud services for scalability, combining LLMs with traditional collaborative filtering models, and regularly evaluating performance metrics ensure the system remains efficient and cost-effective.
Summary.
Answers to Research Questions for Experiment 1 and 2 Answer to RQ1: The incorporation of various goal-oriented prompts impacts the accuracy of GPT-based models, with some prompts enhancing performance while others, like diversity-focused prompts, reducing it. However, GPT-based models generally lag behind CF baselines in zero-shot settings. Answer to RQ2: GPT-based recommendation systems show consistent personalization performance across multiple runs, with minor variability. This stability is comparable to CF baselines, indicating reliable behavior despite inherent randomness in individual runs. Answer to RQ3: GPT-based recommendation systems exhibit lower provider fairness compared to CF baselines. However, incorporating a “Fair Recommender” system role in GPT-based models improves fairness and diversity, demonstrating controllability and potential for mitigation of item unfairness. Answer to RQ4: GPT-based models tend to recommend newer movies and show a broader genre distribution compared to CF baselines, which prefer older, well-established films. Including explicit user ratings in prompts enhances personalization performance significantly. Answer to RQ5: Different ICL strategies impact RecLLM quality and biases variably, with zero-shot learning generally providing better accuracy, while ICL strategies, especially with demographic information, offer improvements in certain contexts. Prompt design plays a crucial role in these outcomes. Answer to RQ6: The economic and practical implications of using GPT-based models include significant inference costs and latency issues. Mitigation strategies like token optimization, batch processing, and using distilled models can help manage these challenges, ensuring efficient and cost-effective system performance.
@@ 6. Conclusion and Future Directions
This research provides a comprehensive analysis of biases of Recommender Systems using Large Language Models (RecLLMs), specifically focusing on ChatGPT-based systems. Our study highlights the distinct capabilities and biases of these systems compared to traditional Collaborative Filtering (CF) methods within the domain of movie and music recommendations. The experimental results emphasize the significant impact of prompt design strategies on various aspects of recommendation quality, including accuracy, provider fairness, catalog coverage, diversity, temporal stability, genre dominance, and recency.
In particular, we conducted two experiments using different datasets to evaluate the effectiveness of these strategies: classical top-$K$ recommendations and sequential in-context learning (ICL).
Key contributions and findings of our work include:
  • Enhanced Fairness and Diversity through Role-Based Prompts: Utilizing role-based prompts (e.g., “act as a recommender” or “act as a fair recommender”) significantly enhances fairness and diversity in recommendations, effectively mitigating item unfairness.
  • Sequential In-Context Learning (ICL): Incorporating sequential in-context learning prompts (ICL-1 and ICL-2) shows improvements in recommendation accuracy in certain contexts, particularly with demographic information. For instance, in the MovieLens dataset with age-group context, the Hit Rate (HR) for ICL-2 improves compared to zero-shot learning. However, ICL prompts do not consistently lead to better item fairness and coverage compared to zero-shot learning.
  • Improved Personalization with Ratings: Incorporating user ratings into prompts improves the personalization of recommendations, highlighting the potential for fine-tuning prompts to better reflect user preferences.
  • Temporal Bias towards Recent Releases: GPT-based models exhibit a strong tendency to recommend newer movies, particularly those released post-2000, contrasting with the older preferences of CF models. This suggests that GPT models can capture recent trends better, making them more temporally relevant.
  • Item Fairness and Diversity: While GPT-based models generally exhibit lower item fairness compared to classical CF models, scenarios involving fair recommender roles show improved fairness and diversity, as indicated by higher entropy values.
  • Stability in Performance Metrics: Despite inherent randomness, GPT-based models show consistent performance across multiple runs, as evidenced by low standard deviations in metrics such as NDCG. This stability is crucial for reliable recommendation systems.
Our research identifies several key areas for further exploration that can push forward the capabilities and effectiveness of recommendation systems leveraging large language models (RecLLMs). Expanding research to include various content domains such as books, music, and e-commerce, along with conducting long-term user engagement studies, will provide deeper insights into this research. Firstly, refining prompt design and enhancing generalization across various datasets and tasks is a crucial step to improving recommendation accuracy and relevance. Additionally, implementing more advanced few-shot learning techniques, such as carefully selecting which examples to use and how to present them, can enhance the contextual relevance and accuracy of recommendations. These examples could be chosen from similar users (peer users) to make the system behave more like CFg models, thereby combining the strengths of both approaches.
Moreover, integrating GPT with collaborative filtering (CF) in a Retrieval-Augmented Generation (RAG) framework could leverage the strengths of both methodologies, leading to more robust and effective recommender systems. Future work should also address additional fairness dimensions, such as ensuring consumer-side fairness to prevent recommendations from disproportionately favoring certain user groups. Calibration, or aligning recommendations with user expectations to offer a balanced mix of familiar and novel items, is another critical aspect. Furthermore, ensuring counterfactual fairness, where recommendations remain fair even when user attributes or behaviors are altered, is also an essential and interesting future direction.
We explicitly acknowledge the limitations of our current work in using a single LLM based on ChatGPT. Future studies should explore other large language models (LLMs) and compare their performance to further understand their potential and limitations in diverse recommendation scenarios.